OpenStudy (anonymous):

a certain number of cats eats a certain number of mice over a period of several days.each cat eats the same number of mice .which is greater than the number of cats.if there are atleast 3 cats and the total number of mice eaten is 999919,how many cats were there?

OpenStudy (anonymous):

say wht?

OpenStudy (anonymous):

Sounds like they want you to factor 999919.

ganeshie8 (ganeshie8):

9 cats

OpenStudy (shubhamsrg):

let c and m be cats and mice now c>=3 if each cat eats n mice, nc=999919=1000000 - 81= 1000^2 - 9^2 = (1009)*(991) also, m>c so c should be the lesser one..991 might be the ans..

OpenStudy (shubhamsrg):

i dont understand what use do we make of this : c>=3

OpenStudy (shubhamsrg):

ohh,,well maybe because c=1 had been another ans of this info was not given..hmmn..

ganeshie8 (ganeshie8):

yh.. @shubhamsrg :) i used the same logic. did a mistake in numbers and got 9

OpenStudy (anonymous):

Yes, that sounds correct. Since the factors are 1,991,1009,999919, they stated there was one than more cat so that someone couldn't say 1 cat ate them all.

OpenStudy (lgbasallote):

isn't this like 3 roosters lay 3 eggs per week. how many eggs will 6 roosters lay? i find it similar ;)